OpenStudy (pflori1234_pop):

f(x)= 3x^2 +5x-z g(x)=5x^2 -3x+z m(x)=2x^2-2x

OpenStudy (pflori1234_pop):

Write an expression to represent 2( f(x) -g(x) ) + 3(m(x))

OpenStudy (pflori1234_pop):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

first tell me what f(x)-g(x) is equal to

OpenStudy (pflori1234_pop):

-2x^2 +7x?

OpenStudy (pflori1234_pop):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

close but no

jimthompson5910 (jim_thompson5910):

5 minus -3 is not 7

OpenStudy (pflori1234_pop):

isn't it 5 +3 cuz the -3 became a positive once i distributed the negative signs to all numbers?

jimthompson5910 (jim_thompson5910):

yes 5 minus -3 = 5+3 = 8

OpenStudy (pflori1234_pop):

ohh oh I'm sorry i dont know what i was thinking

jimthompson5910 (jim_thompson5910):

f(x)-g(x) = -2x^2 + 8x

OpenStudy (pflori1234_pop):

yea

jimthompson5910 (jim_thompson5910):

what is 2*(f(x) - g(x)) equal to?

OpenStudy (pflori1234_pop):

-4x^2 +16x?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

m(x)=2x^2-2x what is 3*m(x) equal to?

OpenStudy (pflori1234_pop):

6x^2 -6x

jimthompson5910 (jim_thompson5910):

2( f(x) -g(x) ) + 3(m(x)) turns into -4x^2+16x + 6x^2-6x simplify that further

OpenStudy (pflori1234_pop):

2x^2 +10?

jimthompson5910 (jim_thompson5910):

close, it's actually 2x^2+10x

OpenStudy (pflori1234_pop):

oh yea i forgot the x

OpenStudy (pflori1234_pop):

thank you!

jimthompson5910 (jim_thompson5910):

sure thing